OpenStudy (anonymous):

PLEASE HELP WILL FAN AND GIVE MEDAL FOR HELPING! Mrs. Walters drove 360 miles on 25 gallons of gas. Mr. Walters drove 480 miles on 32 gallons of gas. Find the mileage per gallon for each car. Which car is getting the better mileage per gallon?

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

well mrs.walters drove 360 miles on 25 gallons of gas so she gets 14.4 mpg and mr walters drove 480 on 32 gallons of gas so he gets 15 mpg so mr.walters is getting the better millage. hope that helps:)

OpenStudy (c1ndrella78):

i am pretty sure that you would divide so 360 divide 25= 480 divide 32= the higher one has better gas mileage

OpenStudy (vijeya3):

Mrs Waters=360/25 miles/gallon that is 14.4miles/gallon Mr Waters=480/32 miles/gallon =15miles/gallon Mr Waters travels more distance in one gallon,so Mr. Waters' s car gives better mileage. If this is helpful,fan and medal me...